I need to prove that the vectors u,v and w are in the span(u,v,w). I don't really know where to start. Any help would be appreciated. Thanks!
Answer by khwang(438)

span(u,v,w) means the linear combinations of the three vectors u,v &w.
Hence, span(u,v,w) = {au+bv+cw| where a,b,c are scalars}
Since u = 1* u + 0*v + 0*w,
v = 0* u + 1*v + 0*w,
and w = 0* u + 0*v + 1*w,
we conclude that u,v and w belongs to span(u,v,w).
This question is pretty basic so you have to work hard.
Kenny